Tyler has a rectangular garden that measures 10 m wide by 13 m long. He wants to increase the area to 208 m2 by increasing the width and length by the same amount. What will be the width (shorter dimension) of the new garden?
Answer by checkley77(12844)

(10+x)(13+x)=208
130+23x+x^2=208
x^2+23x+130-208=0
x^2+23x-78=0
(x+26)(x-3)=0
x-3=0
x=3 ans. for the additional increase of the length & width.
Proof:
(10+3)(13+3)=208
13*16=208
208=208